OpenStudy (anonymous):

Simplify the expression 5 - 6p + 3p - 8 A. -1- p B. -p C. 8- 14p D. 2+ 7p

OpenStudy (midhun.madhu1987):

Combine the like terms like terms...and simplify

OpenStudy (anonymous):

I combined them but I'm stuck on the problem...

OpenStudy (math&ing001):

Something is missing the answer is not on the choices.

OpenStudy (midhun.madhu1987):

can u check the question as well as answers?

OpenStudy (anonymous):

I forgot that it is divided by 3

OpenStudy (math&ing001):

Ah that's the missing part.

OpenStudy (midhun.madhu1987):

:)

OpenStudy (math&ing001):

-6p +3p = ?

OpenStudy (math&ing001):

You're supposed to answer that @Shakita79595

OpenStudy (anonymous):

Is it -9p

OpenStudy (math&ing001):

Nope, -6p + 3p = (-6+3)p = -3p

OpenStudy (anonymous):

Oh ohkay

OpenStudy (math&ing001):

So -3p + 5 - 8 = ?

OpenStudy (math&ing001):

@Shakita79595

OpenStudy (anonymous):

Uhmm -6

OpenStudy (math&ing001):

I have no idea how you got that !

OpenStudy (anonymous):

I add -3p + 5 and subtracted it from 8

OpenStudy (math&ing001):

'p' is a variable, you can't add it to numbers !

OpenStudy (anonymous):

Ohkayye

OpenStudy (math&ing001):

So back to -3p + 5 - 8 = ?

OpenStudy (math&ing001):

Just add the numbers

OpenStudy (anonymous):

-3+5+8 = 10

OpenStudy (math&ing001):

No just (5 - 8), you cannot add the -3p since it's a variable.

OpenStudy (anonymous):

So it's gone be -3

OpenStudy (anonymous):

Or -13

OpenStudy (math&ing001):

Yep -3 so you -3p-3, now divide by 3.